# The generative quantum eigensolver (GQE) and its application for ground state search

### v1

@@ Abstract
We introduce the generative quantum eigensolver (GQE), a novel method for applying classical generative models for quantum simulation. The GQE algorithm optimizes a classical generative model to produce quantum circuits with desired properties. Here, we develop a transformer-based implementation, which we name the generative pre-trained transformer-based (GPT) quantum eigensolver (GPT-QE), leveraging both pre-training on existing datasets and training without any prior knowledge. We demonstrate the effectiveness of training and pre-training GPT-QE in the search for ground states of electronic structure Hamiltonians. GQE strategies can extend beyond the problem of Hamiltonian simulation into other application areas of quantum computing.
@@ 1 Introduction
The field of quantum computing has experienced a remarkable surge, characterized by rapid advancements in the development of quantum devices. Notably, recent research reports the experimental realization of quantum computing with 48 logical qubits [1], marking the onset of the early fault-tolerant quantum computing regime. However, despite these advancements, this regime’s operational number of gates remains limited. Consequently, it is still unclear how these hardware leaps can be effectively translated into practical advantages in the coming decades.
A decade has passed since some of us introduced the variational quantum eigensolver (VQE) [2], which arguably marked a pivotal moment in the field of quantum computing. In VQE, a cost function is minimized by optimizing parameters embedded in a quantum circuit. The variational nature of the algorithm facilitates reducing the circuit depth so they can be implemented on near-term devices. Since its introduction, many quantum algorithms employing variational techniques (variational quantum algorithms: VQA) have been proposed [3, 4]. However, it has been demonstrated that VQAs encounter several issues, particularly with regards to their trainability for large problem instances [5, 6]. This limitation hinders their competitiveness against classical computers when dealing with problems above a certain size. In this work, we aim to circumvent these shortcomings by constructing an orthogonal, rather than complementary, approach to VQAs.
@@
During the same tumultuous decade, modern machine-learning techniques with deep neural networks have revolutionized numerous areas. In particular, there has been significant advancement in generative models for natural language processing. The advent of the Generative Pre-trained Transformer (GPT) [7] marks a milestone in the evolution of artificial intelligence. Forming the basis of Large Language Models (LLMs), GPT-like transformer models have demonstrated exceptional capabilities in understanding and generating human language. Through the simplicity and inherent efficiency of the attention mechanism [8], transformer models have demonstrated extraordinary performance across a wide array of tasks, showcasing their flexibility and expressivity in a variety of domains (e.g., [7, 8, 9, 10]). Recent achievements, highlighted by models like Chinchilla [11], demonstrate how scaling laws in machine learning can inform the efficient allocation of model size for optimized performance, hinting at even greater potential.
@@
Figure 2: Quantum circuit generation in GPT-QE (GQE, which employs a transformer) is explored. We also show the analogy between document generation in Large Language Models (LLMs) and quantum circuit generation in GPT-QE. The details of quantum circuit generation are described in Section 2.2.
@@
Given those significant achievements in classical generative models, incorporating them into quantum computing algorithms could be a pivotal step in overcoming the enduring challenges faced in practical quantum computing applications. Therefore, we propose the generative quantum eigensolver (GQE), which takes advantage of classical generative models for quantum circuits. Specifically, we employ a classical generative model —denoted as $p_{\vec{\theta}}(U)$, with $\vec{\theta}$ as parameters and $U$ as a unitary operator—to define the probability distribution for generating quantum circuits. In simpler terms, we sample quantum circuits according to this distribution. We train $p_{\vec{\theta}}(U)$ so that generated quantum circuits are likely to have desirable properties. We emphasize that, unlike VQA and its variants, no parameters are embedded in the quantum circuit in GQE; notably, and important to their scalability, all the optimizable parameters are in the classical generative model (Fig. 1). We note that some previous works utilize a generative model for generating parameters in VQE [12, 13, 14], but in GQE, the whole circuit structure is determined by a generative model.
@@
In designing the generative model for quantum circuits, we focus on the transformer architecture [8], which achieves significant success as the backbone of large language models. We can describe GQE with a transformer by using an analogy between natural language documents and quantum circuits (Fig. 2). For a given operator pool, defined by a set of unitary operations $\{U_{j}\}_{j=1}^{L}$ (vocabulary), the transformer generates the sequence of indices $j_{1}\dots j_{N}$ corresponding to the unitary operations $U_{j_{1}}\dots U_{j_{N}}$ (words) and constructs the quantum circuit $U_{N}(\vec{j})=U_{j_{N}}\dots U_{j_{1}}$ (document). The rules for generating indices (grammar) are trained so that a cost value calculated by quantum devices decreases. GQE with a transformer is also able to be pre-trained. If we have a dataset given as pairs of index sequences and cost values: $\{\vec{j}_{m},C(\vec{j}_{m})\}_{m=1}^{M}$ (document dataset), we can pre-train the transformer without running quantum devices, as shown in Section 2. Hence, we give the GQE with transformer the name of generative pre-trained transformer-based quantum eigensolver (GPT-QE).
We expect three advantages in training with a transformer instead of a parameterized quantum circuit: ease of optimization, quantum resource efficiency, and customizability. We have been witnessing the impressive optimizability of deep neural networks (DNN) [15, 16, 17] and in this work, all the developments in this field are readily applicable for quantum computing. By using the cost function landscape of DNNs, GPT-QE is potentially unaffected by the core optimization issues of several VQAs. Regarding quantum resource efficiency, we show in Section 2 that the number of quantum circuits run for each step of the optimization in GPT-QE does not explicitly depend neither on the number of parameters nor on the size of the operator pool since it replaces quantum gradient evaluation with sampling and backpropagation. From this feature, we expect to significantly reduce the number of quantum circuits runs compared to conventional VQAs. Additionally, pre-training also constrains the process of running quantum devices for dataset generation, as we note above. As for the customizability of our appraoch, we can append additional conditioned input –e.g., domain knowledge– to the transformer.
To demonstrate a proof of concept of the GPT-QE approach, which can be readily adapted to several families of VQAs, this paper focuses on the ground state search problem. We will explore these other applications of the method in future work. Accurate molecular electronic ground states have significant utility in applications as diverse as drug discovery [18, 19], materials science [20], and environmental solutions [21]. These ground states enable precise simulations of complex molecular structures, accelerating drug development by identifying effective candidates and properties. In materials science, they aid in designing tailored materials [22], from superconductors to catalysts [23]. Additionally, they address environmental challenges by improving energy solutions and enhancing our understanding of relevant chemical processes [24]. It must be acknowledged that ground state search by time-independent means belongs to the complexity class QMA, and it has become uncertain whether this task is a feasible candidate for finding quantum advantage [25]. However, the potential benefits of conceptual understanding and practical applications continue to drive research and development in this field [26].
@@
We note that previous literature [27, 28, 29] proposes methods for training the structure of quantum circuits using machine learning, especially reinforcement learning. Reinforcement learning approaches tend to require a large number of intermediate quantum states in the circuit to determine each action (the next quantum gate to be generated), which leads to an increase in the number of required measurements as the number of gates increases. The Adaptive Derivative-Assembled Pseudo-Trotter ansatz VQE (ADAPT-VQE) [30] also provides a method to adaptively construct the ansatz structure. The method requires running VQE, hence many measurements, to determine the gates to be added as in the case of the reinforcement learning approaches. Conversely, GPT-QE does not require any intermediate measurements, thus potentially significantly reducing the measurement cost when running the algorithm.
@@
The rest of the paper is organized as follows. In Section 2, we describe the details of GQE. Particularly, we construct GPT-QE and describe its training and pre-training scheme. Section 3 is dedicated to demonstrating the performance of the training and pre-training schemes by using the ground state search for the electronic structure Hamiltonians. In Section 4, we summarize what we know of the algorithm so far and suggest future directions of exploration.
@@ 2.1 Generative Quantum Eigensolver
The generative quantum eigensolver (GQE) is an algorithm to search the ground state of a given Hamiltonian $\hat{H}$. Particularly, we focus on the electronic structure problem, where the Hamiltonian is written as the weighted sum of the tensor products of Pauli operators $\hat{P}_{\ell}$: $\hat{H}=\sum_{\ell}h_{\ell}\hat{P}_{\ell}$. To construct the approach of the GQE, we first illustrate our formulation of the generative model of quantum circuits.
We prepare the operator pool $\mathcal{G}=\{U_{j}\}_{j=1}^{L}$, where $U_{j}$ is a unitary operator and $L$ is the size of the operator pool. One of the choices for the operator pool is a set of time evolution operators: $\{e^{i\hat{P}_{j}t_{j}}\}_{j=1}^{L}$, which we use in our numerical experiment. Given a sequence length $N$, we sample the sequence $\vec{j}=\{j_{1},\ldots,j_{N}\}$ according to the parameterized probability distribution $p_{N}(\vec{\theta},\vec{j})$, where $\vec{\theta}=\{\theta_{p}\}_{p=1}^{P}$ are optimizable parameters. Using the sequence $\vec{j}$, we construct the quantum circuit $U_{N}(\vec{j})=U_{j_{N}}\cdots U_{j_{1}}$. We call $p_{N}(\vec{\theta},\vec{j})$ the generative model of quantum circuits. In the rest of the paper, we omit the variable $\vec{\theta}$ for simplicity. The process of sampling the sequence $\vec{j}$ according to $p_{N}(\vec{j})$ and constructing the quantum circuit $U(\vec{j})$ is simply referred to as “sampling the quantum circuit $U(\vec{j})$ according to $p_{N}(\vec{j})$”.
We construct GQE to search for the ground state of $\hat{H}$ with the generative model of quantum circuits $p_{N}(\vec{j})$. The objective of the problem we target is finding $\vec{j}^{\ast}:=\operatorname*{arg\,min}_{\vec{j}}E_{N}(\hat{H},\vec{j})$ with
$E_{N}(\hat{H},\vec{j}):={\textrm{Tr}}\left(\hat{H}U_{N}(\vec{j})\rho_{\textrm{0}}U_{N}(\vec{j})^{\dagger}\right),$ (1)
@@
We note that we need to select the operator pool $\mathcal{G}$ to be expressive enough, so that $E_{N}(\vec{j}^{\ast})$ is close enough to the ground state energy. It should also be noted that we can select $\mathcal{G}$ to accommodate the native operations and topology of each quantum device. We illustrate the specific choice for the operator pool in our numerical experiment. We train $p_{N}(\vec{j})$ so that the generated quantum circuit $U_{N}(\vec{j})$ is likely to produce a low energy quantum state. We call the approach to optimize $p_{N}(\vec{j})$ as the generative quantum eigensolver.
We now emphasize the difference between VQE and GQE. As shown in Fig. 1, in VQE, we embed parameters in the quantum circuit and optimize them to minimize the energy associated with the generated quantum state. In contrast, all parameters in GQE are embedded in the generative model $p_{N}(\vec{j})$. Consequently, the cost function landscapes in GQE and VQE are different; considering the success of training large models with DNN [15, 16, 17], we expect that GQE potentially addresses the issue of trainability in VQE by exploiting the different landscape, which has been now moved on to the classical computer.
An advantage of the GQE approach is that we are free to choose the generative model $p_{N}(\vec{j})$ from a very rich potential set of families of generative models stemming from the field of machine learning, such as autoencoders, generative adversarial networks, diffusion models, flow models, etc. This paper focuses on the model where the Transformer implements the generative model [8], which achieves significant success as the cornerstone of large language models. In the following subsection, we describe the details of GQE implemented in this transformer setting.
@@ 2.2 GPT Quantum Eigensolver
We construct the specific GQE algorithm using the transformer architecture and provide its training scheme. As we will show later, the approach also involves pre-training; therefore, we call the method generative pre-trained transformer-based quantum eigensolver (GPT-QE). In the following, we describe how the transformer generates quantum circuits. Then, we construct its training/pre-training scheme of GPT-QE.
Quantum circuits generation in GPT-QE
@@
The sequence generation using the transformer can be written as the repetitions of (i) calculating the logit (logarithmic probability) with which each token is generated and (ii) sampling a token according to the corresponding probability distribution. We write the function for the probability calculation as GPT and the function for sampling as Sample.
The function GPT takes the variable-length inputs $\vec{j}^{(k)}=\{j_{1},\ldots,j_{k}\}$, where each element takes an integer value between $1$ and $L$ (the size of the operator pool $\mathcal{G}$). Then it outputs the sequence of logits $W^{(k)}=\{\vec{w}^{(1)},\ldots,\vec{w}^{(k)}\}$ with the same length, where each logit $\vec{w}^{(r)}$ is a real vector of size $L$. We note that GPT has optimizable parameters $\vec{\theta}$ included in the transformer’s architecture, which is not explicitly written in the notation GPT for simplicity.
The function GPT follows the methodology outlined in [7]. Here, we present a concise overview of GPT’s operation: Initially, GPT converts each token in the input sequence into a unique embedding, represented as a vector. These embeddings are transformed by means of multiple attention layers. Each attention layer takes a sequence of vectors as input, with the first layer’s input being the embeddings themselves. The output of each attention layer is also a sequence of vectors, maintaining the same sequence length and vector dimension as the input. This consistency allows the output of one layer to serve as the input for the subsequent layer. The final attention layer’s output is converted into a sequence of logits, constituting the output of GPT. Given an attention layer’s input $\{\vec{v}_{1},\cdots,\vec{v}_{k}\}$, the output can be expressed as $\{\vec{a}_{1},\cdots,\vec{a}_{k}\}$, where each $\vec{a}_{r}$ represents the attention with the same dimension as $\vec{v}_{r}$. The attention $\vec{a}_{r}$ encapsulates the relationship between $\vec{v}_{r}$ and other elements of the input. Notably, through the process of causal masking, $\vec{a}_{r}$ depends solely on preceding elements, i.e., $\{\vec{v}_{r^{\prime}}\}_{r^{\prime}<r}$. The standard implementation of attention computation involves running multiple attention mechanisms in parallel, and their outputs are combined into $\{\vec{a}_{1},\cdots,\vec{a}_{k}\}$ through a weighted average (multi-head attention). For a more comprehensive explanation, see [7].
@@
The function Sample is a stochastic function that takes a logit $\vec{w}=\{w_{j}\}_{j=1}^{L}$ as its input and returns one of the tokens $j\in\{1,\ldots,L\}$. The probability that the token $j$ is sampled is proportional to $e^{-\beta w_{j}}$, with $\beta>0$ a hyper-parameter we can choose. It is customary to sample from the logits, which can be understood as each $j$ being sampled according to the energy $w_{j}$ and the inverse temperature $\beta$ as in statistical mechanics. For simplicity, we omit the variable $\beta$ from the function Sample.
@@
  • In the $k$-th step we sample the token $j_{k}$ with the input $\{0,j_{1},\ldots,j_{k-1}\}$:

### v2

@@ Abstract
[…]
*Abstract We introduce the generative quantum eigensolver (GQE), a new quantum computational framework that operates outside the variational quantum algorithm paradigm by applying classical generative models to quantum simulation. The GQE algorithm optimizes a classical generative model to produce quantum circuits with desired properties. Here, we develop a transformer-based implementation, which we name the generative pre-trained transformer-based (GPT) quantum eigensolver (GPT-QE). We show a proof-of-concept of training and pretraining of GPT-QE applied to electronic structure Hamiltonians, and demonstrate its ability illustrated by surpassing coupled cluster singles and doubles (CCSD) for the strong bond dissociation of the nitrogen molecule and approaching chemical accuracy. We also demonstrate the method on real quantum hardware.
@@ 1 Introduction
The field of quantum computing has experienced a remarkable surge, characterized by rapid advancements in the development of quantum devices. Notably, recent publication reports the experimental realization of quantum computing with 48 logical qubits [1], marking the onset of the early fault-tolerant quantum computing regime. However, despite these advancements, this regime’s operational number of gates remains limited. Consequently, it is still unclear how these hardware leaps can be effectively translated into practical advantages in the coming years.
A decade has passed since some of us introduced the variational quantum eigensolver (VQE) [2], which arguably marked a pivotal moment in the field of quantum computing. In VQE, a cost function is minimized by optimizing parameters embedded in a quantum circuit. The variational nature of the algorithm facilitates reducing the depth of the circuit so they can be implemented on near-term devices. Since its introduction, many quantum algorithms employing variational techniques (variational quantum algorithms: VQAs) have been proposed [3, 4]. However, it has been demonstrated that VQAs encounter several issues, particularly with regards to their trainability for large problem instances [5, 6]. This limitation hinders their competitiveness against classical computers when dealing with problems above a certain size. In this work, we aim to circumvent these shortcomings by constructing a completely different approach that operates outside the VQA paradigm.
@@
During the same tumultuous decade, modern machine-learning techniques based on deep neural networks have revolutionized numerous areas. In particular, there has been significant advancement in generative models for natural language processing. The advent of the Generative Pre-trained Transformer (GPT) [7] marks a milestone in the evolution of artificial intelligence. Forming the basis of Large Language Models (LLMs), GPT-like transformer models have demonstrated exceptional capabilities in understanding and generating human language. Through the simplicity and inherent efficiency of the attention mechanism [8], transformer models have demonstrated extraordinary performance across a wide array of tasks, showcasing their flexibility and expressivity in a variety of domains (e.g., [7, 8, 9, 10]). Recent achievements, highlighted by models like Chinchilla [11], demonstrate how scaling laws in machine learning can inform the efficient allocation of model size for optimized performance, hinting at even greater potential.
@@
Figure 2: Depiction of quantum circuit generation in GPT-QE (GQE, which employs a transformer). We also show the analogy between document generation in Large Language Models (LLMs) and quantum circuit generation in GPT-QE. The details of quantum circuit generation are described in Section 2.2.
@@
Given those significant achievements in classical generative models, incorporating generative models into quantum computing algorithms could be a pivotal step in overcoming the enduring challenges faced in practical quantum computing applications. Therefore, we propose the generative quantum eigensolver (GQE), which takes advantage of classical generative models for quantum circuits. Specifically, we employ a classical generative model —denoted as $p_{\vec{\theta}}(U)$, with $\vec{\theta}$ as parameters and $U$ as a unitary operator—to define a probability distribution for generating quantum circuits. In simpler terms, we sample quantum circuits according to this distribution. We train $p_{\vec{\theta}}(U)$ so that generated quantum circuits are likely to have desirable properties. We emphasize that, unlike VQA and its variants, no parameters are embedded in the quantum circuit in GQE; notably, and importantly for scalability, all the optimizable parameters are in the classical generative model (Fig. 1). We note that some previous works utilize a generative model for generating parameters in VQE [12, 13, 14], but in GQE, the whole circuit structure is determined by a generative model.
@@
In designing the generative model for quantum circuits generation, we focus on the transformer architecture [8], which has achieved significant success as the backbone of large language models. We can describe the implementation of GQE with a transformer by using an analogy between natural language documents and quantum circuits (Fig. 2). For a given operator pool, defined by a set of unitary operations $\{U_{j}\}_{j=1}^{L}$ (vocabulary), the transformer generates the sequence of indices $j_{1}\dots j_{N}$ corresponding to the unitary operations $U_{j_{1}}\dots U_{j_{N}}$ (words) and constructs the quantum circuit $U_{N}(\vec{j})=U_{j_{N}}\dots U_{j_{1}}$ (document). The rules for generating indices (grammar) are trained so that a cost value calculated by quantum devices decreases. GQE with a transformer is also able to be pre-trained. If we have a dataset given as pairs of index sequences and cost values: $\{\vec{j}_{m},C(\vec{j}_{m})\}_{m=1}^{M}$ (document dataset), we can pre-train the transformer without running quantum devices, as shown in Section 2. Hence, we give the GQE with transformer the name of generative pre-trained transformer-based quantum eigensolver (GPT-QE).
To demonstrate the effectiveness of our fundamentally new GPT-QE approach, this paper demonstrates solving the molecular electronic ground state search problem, which has significant utility in applications including drug discovery [15, 16] and materials science [17]. Our results demonstrate that GPT-QE outperforms coupled cluster singles and doubles (CCSD) methods in the strong bond dissociation regime of the nitrogen molecule.
An important distinction from traditional VQE approaches is the potential for incorporating conditional inputs into the generative model. While this paper does not directly demonstrate this capability, the transformer architecture enables the integration of domain knowledge or problem-specific constraints as conditional inputs during circuit generation. This represents a fundamentally different paradigm compared to VQE, where the circuit structure is typically fixed and only parameters are optimized. We discuss this capability and its implications in Section 2.2. Following the initial preprint of this work, subsequent research has successfully demonstrated this conditional input capability for combinatorial optimization problems [18], validating the potential of this approach.
@@
We note that previous literature [19, 20, 21] proposes methods for constructing the structure of quantum circuits using machine learning, especially reinforcement learning. Reinforcement learning approaches tend to require access to a large number of intermediate quantum states in the circuit to determine each action (the next quantum gate to be generated), which leads to an increase in the number of required measurements as the number of gates increases. Conversely, GPT-QE does not require any intermediate measurements, thus potentially significantly reducing the measurement cost when running the algorithm.
@@
The rest of the paper is organized as follows. In Section 2, we describe the details of GQE. Particularly, we construct GPT-QE and describe its training and pre-training scheme. Section 3 is dedicated to demonstrating the training and pre-training schemes by using the ground state search for the electronic structure Hamiltonians as a benchmark. It also presents the results of GPT-QE on a noisy real device, with particular focus on the effectiveness of error mitigation techniques in improving the training process. In Section 4, we summarize what we know about the algorithm so far and suggest future directions of exploration.
@@ 2.1 Generative Quantum Eigensolver
The generative quantum eigensolver (GQE) is an algorithm to search for the ground state of a given Hamiltonian $\hat{H}$. Particularly, we focus on the electronic structure problem, where the Hamiltonian is written as the weighted sum of the tensor products of Pauli operators $\hat{P}_{\ell}$: $\hat{H}=\sum_{\ell}h_{\ell}\hat{P}_{\ell}$. To construct the approach of GQE, we first illustrate our formulation of the generative model of quantum circuits.
We prepare the operator pool $\mathcal{G}=\{U_{j}\}_{j=1}^{L}$, where $U_{j}$ is a unitary operator and $L$ is the size of the operator pool. One of the possible choices for the operator pool is a set of time evolution operators: $\{e^{i\hat{P}_{j}t_{j}}\}_{j=1}^{L}$, which is what we use in our numerical experiments. The detailed configuration of our operator pool is described in Appendix A. Given a sequence length $N$, we sample the sequence $\vec{j}=\{j_{1},\ldots,j_{N}\}$ according to the parameterized probability distribution $p_{N}(\vec{\theta},\vec{j})$, where $\vec{\theta}=\{\theta_{p}\}_{p=1}^{P}$ are optimizable parameters. Using the sequence $\vec{j}$, we construct the quantum circuit $U_{N}(\vec{j})=U_{j_{N}}\cdots U_{j_{1}}$. We call $p_{N}(\vec{\theta},\vec{j})$ the generative model of quantum circuits. In the rest of the paper, we omit the variable $\vec{\theta}$ for simplicity. The process of sampling the sequence $\vec{j}$ according to $p_{N}(\vec{j})$ and constructing the quantum circuit $U(\vec{j})$ is simply referred to as “sampling the quantum circuit $U(\vec{j})$ according to $p_{N}(\vec{j})$”.
We construct GQE to search for the ground state of $\hat{H}$ using the generative model of quantum circuits $p_{N}(\vec{j})$. The objective of the problem we target is finding $\vec{j}^{\ast}:=\operatorname*{arg\,min}_{\vec{j}}E(\hat{H},\vec{j})$ with
$E(\hat{H},\vec{j}):={\textrm{Tr}}\left(\hat{H}U_{N}(\vec{j})\rho_{\textrm{0}}U_{N}(\vec{j})^{\dagger}\right),$ (1)
@@
We note that we need to select the operator pool $\mathcal{G}$ to be expressive enough, so that $E(\vec{j}^{\ast})$ is close enough to the ground state energy. It should also be noted that we can select $\mathcal{G}$ to accommodate the native operations and topology of each quantum device, or to reflect domain-specific insights. We illustrate a specific choice for a chemistry-inspired operator pool in our numerical experiment. We train $p_{N}(\vec{j})$ so that the generated quantum circuits $U_{N}(\vec{j})$ are likely to produce a low energy quantum state. We call the approach to optimize $p_{N}(\vec{j})$ as the generative quantum eigensolver.
We now emphasize the difference between VQE and GQE. As shown in Fig. 1, in VQE, we embed parameters in the quantum circuit and optimize them to minimize the energy associated with the generated quantum state. In contrast, all parameters in GQE are embedded in the generative model $p_{N}(\vec{j})$. Consequently, the cost function landscapes in GQE and VQE are different; considering the success of training large models with DNN [22, 23, 24], we expect that GQE can potentially address certain issues of trainability in VQE by exploiting the different landscape, which has been now moved onto the classical computer.
An advantage of the GQE approach is that we are free to choose the generative model $p_{N}(\vec{j})$ from a very rich potential set of families of generative models stemming from the field of machine learning. Another advantage is the ability to incorporate contextual inputs, enabling the generation of problem-specific quantum circuits based on classical information such as molecular geometry or optimization constraints, as demonstrated in combinatorial optimization applications [18].
@@ 2.2 GPT Quantum Eigensolver
We construct a specific GQE algorithm using the transformer architecture and provide its training scheme. As we will show later, the approach also involves pre-training; therefore, we call the method Generative pre-trained transformer-based quantum eigensolver (GPT-QE). In the following, we describe how the transformer generates quantum circuits. Then, we construct the training/pre-training scheme of GPT-QE.
2.2.1 Quantum circuits generation in GPT-QE
@@
The sequence generation using the transformer can be written as repetitions of (i) calculating the logit (logarithmic probability) with which each token is generated and (ii) sampling a token according to the corresponding probability distribution. We write the function for the logit calculation as GPT and the function for sampling as Sample.
The function GPT takes the variable-length inputs $\vec{j}^{(k)}=\{j_{1},\ldots,j_{k}\}$, where each element takes an integer value between $1$ and $L$ (the size of the operator pool $\mathcal{G}$). Then it outputs a sequence of logits $W^{(k)}=\{\vec{w}^{(1)},\ldots,\vec{w}^{(k)}\}$ with the same length, where each logit $\vec{w}^{(r)}$ is a real vector of size $L$. We note that GPT has optimizable parameters $\vec{\theta}$ included in the transformer’s architecture, which is not explicitly written in the notation GPT for simplicity.
The function GPT follows the methodology outlined in [7]. Here, we present a concise overview of GPT’s operation: Initially, GPT converts each token in the input sequence into a unique embedding, represented as a vector. These embeddings are transformed by means of multiple attention layers. Each attention layer takes a sequence of vectors as input, with the first layer’s input being the embeddings themselves. The output of each attention layer is also a sequence of vectors, maintaining the same sequence length and vector dimension as the input. This consistency allows the output of one layer to serve as the input for the subsequent layer. The final attention layer’s output is converted into a sequence of logits, constituting the output of GPT. Given an attention layer’s input $\{\vec{v}_{1},\cdots,\vec{v}_{k}\}$, the output can be expressed as $\{\vec{a}_{1},\cdots,\vec{a}_{k}\}$, where each $\vec{a}_{r}$ represents the attention with the same dimension as $\vec{v}_{r}$. The attention $\vec{a}_{r}$ encapsulates the relationship between $\vec{v}_{r}$ and other elements of the input. Notably, through the process of causal masking, $\vec{a}_{r}$ depends solely on preceding elements, i.e., $\{\vec{v}_{r^{\prime}}\}_{r^{\prime}<r}$. The standard implementation of attention computation, as employed here, involves running multiple attention mechanisms in parallel, and their outputs are combined into $\{\vec{a}_{1},\cdots,\vec{a}_{k}\}$ through a weighted average (multi-head attention). For a more comprehensive explanation, see [7].
@@
The function Sample is a stochastic function that takes a logit $\vec{w}=\{w_{j}\}_{j=1}^{L}$ as its input and returns one of the tokens $j\in\{1,\ldots,L\}$. The probability that the token $j$ is sampled is proportional to $e^{-\beta w_{j}}$, with $\beta>0$ a hyper-parameter we can choose. This can be understood as each $j$ being sampled according to the energy $w_{j}$ and the inverse temperature $\beta$ as in statistical mechanics. For simplicity, we omit the variable $\beta$ from the function Sample.
@@
  • In the $k$-th step, we sample the token $j_{k}$ with the input $\{0,j_{1},\ldots,j_{k-1}\}$: